# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 4.1 Overview and Comparative Analysis
Prompt-based datasets represent a major class of benchmarks for evaluating fairness in language models (LMs). In contrast to counterfactual datasets that focus on paired sentence comparisons under controlled attribute changes, prompt-based datasets assess fairness by analyzing the model’s generated content given a fixed prompt. This design better reflects real-world LM usage, including text generation, dialogue, and question answering (QA).
@@
BOLD Sentence completion Wikipedia prompts Demographic regard (gender, race, religion, etc.) English
@@
RealToxicityPrompts Open-ended generation Reddit-based prompts (filtered) Toxicity trigger bias English
@@
HONEST Fill-in-the-blank Template-based Gender bias in occupation completions English (+ Fr., Fil., Ger., etc.)
@@
TrustGPT QA / judgment task Templates + model completions Trustworthiness, intersectional bias English
BBQ Multiple-choice QA Manual demographic templates 9 demographic dimensions, ambiguity bias English
UnQover Masked factual QA Handwritten probes Stereotype-triggered factual errors English
Grep-BiasIR QA / IR-style scoring Demographic queries + LM retrieval Retrieval and association bias English
@@
Table 7: Overview of Counterfactual Fairness Datasets. The column headings follow the taxonomy defined in Section 2.2. For the complementary list of counterfactual datasets, see Table 1.
The seven datasets reviewed in this section—BOLD, RealToxicityPrompts, HONEST, TrustGPT, BBQ, UnQover, and Grep-BiasIR—exhibit diverse structural formats, source constructions, linguistic scopes, and bias typologies. We organize their comparison in Table 7, which summarizes key taxonomic dimensions for each dataset.
Structurally, prompt-based datasets fall into two categories: sentence completion and QA-style prompting. Sentence completion datasets such as BOLD, RealToxicityPrompts, and HONEST typically involve open-ended or fill-in-the-blank formats, revealing biases in spontaneous generation, such as stereotype evocation or toxic continuations. In contrast, QA-based datasets—like BBQ, UnQover, TrustGPT, and Grep-BiasIR—test model reasoning under ambiguity or constraint, enabling finer-grained diagnosis of biased inference mechanisms.
The source of prompts plays a significant role in determining lexical style and demographic coverage. Some datasets use naturally occurring prompts from corpora (e.g., Wikipedia in BOLD), while others rely on template-based constructions authored or curated by researchers (e.g., BBQ, HONEST). Each approach has advantages: natural prompts preserve ecological validity, while templates allow precise control over demographic features and semantic variation.
Linguistically, most datasets are English-centric, with HONEST being one of the few to offer multilingual variants. This English dominance reflects broader trends in NLP, but also underscores the need for more cross-linguistic fairness evaluation tools.
Regarding bias typology, datasets vary in their focus and granularity. BBQ has the broadest demographic axis coverage (nine dimensions), while RealToxicityPrompts concentrates on toxicity and implicit stereotypes. BOLD targets sentiment and regard across demographic domains, and HONEST focuses on gender-inflected sentence completions. Grep-BiasIR and TrustGPT include model-in-the-loop retrieval or judgment, expanding evaluation to dynamic LM interactions. Construction biases (e.g., from prompt source or annotation method) are also present to varying degrees and must be accounted for when interpreting results.
The datasets listed in Table 7 illustrate the richness and diversity of prompt-based fairness evaluations. Each structural and methodological design choice—whether in prompt sourcing, task format, or annotation—shapes the types of biases that can be observed and the conclusions that can be drawn. These distinctions motivate the task-specific analyses that follow in Sections 5.2 and 5.3.
4.2 Sentence Completion Datasets
Sentence completion datasets evaluate fairness by assessing the content generated by LMs when provided with an initial text prompt. This format is particularly useful for understanding biases that emerge in open-ended text generation scenarios, such as dialogue systems or creative writing assistance. Evaluations often focus on the propensity of the model to generate toxic language, perpetuate stereotypes, or exhibit other forms of harmful bias in its continuations.
@@ 4.2.1 BOLD (Bias in Open-Ended Language Generation Dataset)
Description. BOLD is a large-scale benchmark dataset specifically designed to evaluate social biases in open-ended English language generation. Its central objective is to provide a systematic framework for analyzing bias across a broad range of demographic attributes. By leveraging naturally occurring language data, BOLD enables researchers to assess how models propagate or amplify social stereotypes during text generation.
Dataset Taxonomy. Structurally, BOLD is categorized as a sentence completion dataset. Prompts were derived from English Wikipedia through a careful curation process, involving both random and targeted sampling strategies. Specific filters—such as content exclusions and length constraints—were applied to ensure representativeness and relevance across the targeted domains. The dataset exclusively covers English-language content and focuses on five core bias domains: profession, gender, race, religion, and political ideology. Within these domains, it captures 43 demographic subgroups. However, due to its reliance on Wikipedia, the dataset may inherit systemic construction biases. The original authors caution against assuming Wikipedia as an unbiased gold standard. BOLD is publicly accessible via GitHub and the Hugging Face Hub, facilitating widespread use and reproducibility.
@@
Intrinsic Characteristics. The dataset comprises 23,679 unique prompts, each constructed by identifying Wikipedia sentences in which a group identifier (e.g., a profession or gendered term) appears early. Typically, the first five words of the sentence are retained and concatenated with the demographic term to form a short prompt, ranging from six to nine words. To mitigate direct metric confounding, proper names and identifiers are anonymized in the evaluation phase. For bias measurement, the creators developed a suite of automated metrics, including toxicity (via fine-tuned BERT), sentiment (using VADER), regard (via pre-trained BERT), gender polarity (via word embeddings and unigram matching), and psycholinguistic norms such as valence, arousal, and basic emotions. These automated evaluations were validated against human judgments obtained through crowdsourcing, enhancing their reliability.
@@
Domain Focus and Significance. BOLD is tailored to assess bias in the five aforementioned demographic domains as represented in Wikipedia. Its design allows for evaluating how generative models respond to demographic cues embedded in realistic, open-ended prompts.
@@
Strengths and Limitations. Among BOLD’s major strengths are its large scale, ecological validity through the use of natural text, and coverage of multiple socially relevant dimensions. Furthermore, it offers a diverse array of metrics, many of which have been cross-validated with human annotation. Nevertheless, its limitations are notable. The dataset is restricted to English and heavily dependent on Wikipedia, which may introduce unaccounted-for societal and cultural biases. Demographic categorization is not exhaustive—gender remains binary, and race is treated largely within a U.S. framework. Moreover, the absence of controlled template structures means that confounding linguistic or contextual factors may be present in the prompts, making it difficult to isolate the bias attributable solely to the model rather than the data source itself.
@@
Representativeness Bias. Following the protocol of Section 2.3 we compare the empirical frequency of each protected attribute in BOLD to U. S. Bureau of Labor Statistics (BLS 2024) population benchmarks. The Kullback–Leibler divergences (Eq. 1) are:
@@
Table 8: KL divergences between BOLD and U.S. population benchmarks (BLS 2024), indicating representativeness bias across demographic domains. Higher KL values denote greater distributional mismatch. The “Imbalance Pattern” column highlights which groups are over- or under-represented in BOLD.
@@
The race domain shows the largest divergence: BOLD prompts are 63.2 % White but only 1.3 % Hispanic, whereas the reference distribution is 57.8 % White and 18.7 % Hispanic. The religion domain over-represents non-Christian topics ($73$% vs. $37$%), reflecting Wikipedia’s emphasis on minority-religion articles. Gender imbalance is comparatively small ($\mathrm{KL}=0.033$) but still skews male (64 % vs. 51 %). These divergences confirm that BOLD inherits Wikipedia’s systemic coverage biases and is not demographically representative of the U.S. population, an important caveat when attributing biases observed in model generations.
Annotation Bias. Under our definition (2.3), annotation bias arises when the labels or scores embedded in the dataset—whether provided by human raters or an automatic instrument—systematically differ across demographic groups in ways unrelated to ground truth. BOLD, however, is distributed as a prompt-only corpus: no gold labels, polarity scores, or human judgments are attached at the item level. The sole metadata field is the deterministic mapping of each prompt to one of 43 demographic sub-groups, which the authors derive from Wikipedia category names via rule-based matching. Consequently, BOLD exhibits no classical annotation bias. Any disparities observed during evaluation stem from the auxiliary metrics (Perspective toxicity, VADER sentiment, etc.) applied post-hoc to generated text; these fall under differential metric bias rather than annotation bias within the dataset itself.
@@
Stereotype Leakage. We probed the co-occurrence structure between a predefined set of group tokens ($|W_{G}|{=}58$) and trait tokens ($|W_{T}|{=}178$). The aggregate mutual information is
@@
indicating that the demographic cue in a prompt decreases uncertainty about the accompanying occupation/descriptor by almost one bit—non-trivial leakage for a corpus whose prompts average only 7 tokens.
Figure 29 ranks the ten strongest group–trait pairings by pointwise mutual information (PMI). Although the top pairs (him–warrior, elderly–psychiatrist, men–builder, …) attain PMI$>7$, eight of them occur $\leq 4$ times, illustrating the well-known “rare-event inflation” of PMI. Nevertheless, these outliers surfaced long-tail stereotypes that template benchmarks often miss (e.g., gentleman–valet, native–interpreter).
@@
Differential Metric Bias.
(a) Gender Polarity. To evaluate gender bias in the BOLD dataset, we examine the gender polarity of generated texts across four demographic domains—political ideology, profession, race, and religion—using three metrics: Unigram Matching, Gender-Max, and Gender-Wavg. These metrics respectively capture direct mentions of gendered words and the projection of semantic content along a gender direction in embedding space. We omit the gender domain from this analysis, as prompts within that domain are explicitly constructed around gender categories (e.g., “American actors” vs. “American actresses”), rendering gender polarity measurements tautological and analytically uninformative.
We illustrate results from the Profession Domain as a representative example. In Figure 30, traditionally male-dominated professions—such as engineering, scientific, and military occupations—are associated with completions containing more male terms. Conversely, professions like nursing, dance, and writing are linked with stronger female-aligned word use. The Gender-Max scores peak near +0.6 for female-oriented professions and drop below -0.4 for technical fields, confirming that gender-stereotypical associations are preserved in completions. Gender-Wavg scores echo this pattern with moderated magnitudes. These findings underline the persistence of occupational gender stereotypes in open-ended generation.
Results for the Political Ideology, Race, and Religion domains are provided in Appendix B.1.
Figure 30: Gender Polarity Distribution in Profession Domain
(b) Regard. We analyze the regard expressed in generated completions, a metric that captures the level of social respect or esteem toward different demographic groups [99]. It classifies text as expressing positive, neutral, or negative regard.
Detailed regard distribution results for Political Ideology, Profession, Race, and Religion domains are included in Appendix B.2.
Figure 31: Regard Distribution and Average Scores in Gender Domain
(c) Sentiment. We compute sentiment using the VADER tool [101], which scores text on a compound scale from $-1$ (negative) to $+1$ (positive).
Sentiment patterns across Political Ideology, Profession, Race, and Religion domains are presented in Appendix B.3.
Figure 32: Sentiment Distribution and Average Scores in Gender Domain
(d) Toxicity. We analyze toxicity using the Perspective API across multiple subdimensions: toxic, severe toxic, identity attack, insult, threat, and obscene.
Toxicity analyses for Gender, Profession, Race, and Religion domains are reported in Appendix B.4.
Figure 33: Toxicity Scores in Political Ideology Domain
In summary, while average toxicity scores in BOLD are generally low, their distribution is not uniform. Certain demographic groups—especially political ideologies like fascism, racial and religious minorities, and atheists—are disproportionately exposed to toxic outputs. These disparities illustrate how language models may perpetuate harmful social stereotypes, and underscore the need for granular toxicity evaluation across multiple labels when assessing bias in generative systems.

### v2

@@ 4.1 Overview and Comparative Analysis
Open-ended evaluation datasets assess fairness by analyzing the content generated by language models in response to given prompts. Unlike constrained-form datasets, which focus on tasks with predefined outputs, open-ended datasets allow models to generate free-form text, making them more reflective of real-world applications such as text generation, dialogue, and creative writing. These datasets examine biases in the model’s output, including toxicity, stereotype reinforcement, and other demographic-related biases.
@@
BOLD (Sec. 4.1.1) Sentence completion Wikipedia-based prompts Bias in sentiment and demographic regard (gender, race, religion) English
@@
RealToxicityPrompts (Sec. 4.1.2) Open-ended generation Reddit-based prompts (filtered for toxicity) Toxicity, implicit stereotypes, offensive language English
@@
HONEST (Sec. 4.1.3) Fill-in-the-blank Template-based Gender bias in occupational role completions (e.g., “doctor”, “nurse”) English (+ French, Filipino, German, etc.)
@@
TrustGPT (Sec. 4.1.4) QA / judgment task Templates + model completions Trustworthiness, intersectional bias (e.g., gender and race in responses) English
@@
Table 7: Overview of Open-ended Evaluation Datasets. The column headings follow the taxonomy defined in Section 2.1. For the complementary list of constrained-form datasets, see Table 1.
The four datasets discussed in this section, BOLD, RealToxicityPrompts, HONEST, and TrustGPT, are designed to uncover various biases in open-ended generation tasks. These datasets are summarized in Table 7, which highlights their structural formats, sources, and bias typologies.
Structurally, these datasets differ in the types of prompts they employ. BOLD and RealToxicityPrompts primarily focus on sentence completion or open-ended text generation tasks, aiming to expose biases such as toxic language and stereotype reinforcement. In contrast, HONEST and TrustGPT feature more complex task designs, including fill-in-the-blank and judgment tasks, which assess model biases in specific domains like gender roles and trustworthiness.
The sources of the prompts also vary. BOLD utilizes naturally occurring prompts sourced from Wikipedia, while HONEST and TrustGPT rely on template-based constructions. This approach allows for more controlled experimentation with demographic variables, such as gender and occupation, which is essential for evaluating how models respond to sensitive content.
While the datasets predominantly focus on English-language prompts, HONEST is a notable exception, offering multilingual support with versions in French and Filipino. This multilingual aspect helps to broaden the scope of fairness evaluation beyond English-centric dominance in NLP research.
In terms of bias typologies, the datasets differ in their focus. BOLD and RealToxicityPrompts focus on detecting toxicity and harmful language in the generated outputs. HONEST addresses gender biases in occupation-related completions, and TrustGPT evaluates fairness in model decision-making, particularly with respect to trustworthiness and intersectional biases.
All four datasets are publicly accessible, though the mechanisms for access vary. BOLD and RealToxicityPrompts are readily available through platforms like Hugging Face and other open repositories, ensuring easy access for the research community. HONEST is hosted on GitHub with detailed instructions for use, while TrustGPT requires users to engage with its API or pre-configured tools to evaluate model biases.
@@ 4.1.1 BOLD (Bias in Open-Ended Language Generation Dataset)
Description. BOLD is a large-scale dataset specifically designed to evaluate social biases in open-ended English-language generation. Its central objective is to provide a systematic framework for analyzing bias across a broad range of demographic attributes. By leveraging naturally occurring language data, BOLD enables researchers to assess how models propagate or amplify social stereotypes during text generation.
Dataset Taxonomy. BOLD is classified as an open-ended evaluation dataset, requiring models to generate text freely in response to prompts without predefined output constraints. It leverages naturally occurring language data from Wikipedia, constructing 23,679 unique prompts by identifying sentences where group identifiers appear early and retaining the first five words concatenated with demographic terms. The dataset employs a comprehensive suite of automated metrics including toxicity, sentiment, regard, gender polarity, and psycholinguistic norms, enabling systematic analysis of bias across five core demographic domains: profession, gender, race, religion, and political ideology.
In terms of source, BOLD employs natural text sources, with prompts derived from English Wikipedia through a careful curation process involving both random and targeted sampling strategies. This approach provides rich linguistic diversity and represents real-world language usage, offering a more comprehensive view of how biases might manifest in actual communication. However, the challenge with natural text sources is that they may already contain societal biases, making it difficult to discern whether bias comes from the model itself or the underlying data.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English-language content. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, BOLD targets multiple demographic characteristic biases across five core domains: profession, gender, race, religion, and political ideology. Within these domains, it captures 43 demographic subgroups, providing comprehensive coverage of protected attributes. However, due to its reliance on Wikipedia, the dataset may inherit systemic construction biases, particularly representative bias from non-representative sampling of Wikipedia content.
Finally, BOLD demonstrates high accessibility as a publicly available dataset via GitHub and the Hugging Face Hub. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Intrinsic Characteristics. The dataset comprises 23,679 unique prompts, each constructed by identifying Wikipedia sentences in which a group identifier (e.g., a profession or gendered term) appears early. Typically, the first five words of the sentence are retained and concatenated with the demographic term to form a short prompt, ranging from six to nine words. To mitigate direct metric confounding, the proper names and identifiers are anonymized in the evaluation phase. For bias measurement, the creators developed a suite of automated metrics, including toxicity (via fine-tuned BERT), sentiment (using VADER), regard (via pre-trained BERT), gender polarity (via word embeddings and unigram matching), and psycholinguistic norms such as valence, arousal, and basic emotions. These automated evaluations were validated against human judgments obtained through crowdsourcing, enhancing their reliability.
@@
Domain Focus and Significance. BOLD is tailored to assess bias in the five aforementioned demographic domains as represented in Wikipedia. Its design allows for evaluating how generative models respond to demographic cues embedded in realistic open-ended prompts.
@@
Strengths and Limitations. Among BOLD’s major strengths are its large scale, ecological validity through the use of natural text, and coverage of multiple socially relevant dimensions. Furthermore, it offers a diverse array of metrics, many of which have been cross-validated with human annotation. Nevertheless, its limitations are notable. The dataset is restricted to English and heavily dependent on Wikipedia, which may introduce unaccounted-for societal and cultural biases. Demographic categorization is not exhaustive, as gender remains binary, and race is treated largely within a U.S. framework. Moreover, the absence of controlled template structures means that confounding linguistic or contextual factors may be present in the prompts, making it difficult to isolate the bias attributable solely to the model rather than the data source itself.
@@
Representativeness Bias. Following the protocol of Section 2.3 we compare the empirical frequency of each protected attribute in BOLD to U.S. Census 2020 (for race and gender), 2022 American Community Survey (for nationality and disability), Pew Research Center 2021 (for religion and income), Gallup 2023 (for sexual orientation) and CDC 2021 BMI statistics (for weight status) population datasets.  [^10] Top-level Wikipedia categories mapped to census classes are shown in the Appendix A.1.2. Gender uses American actors vs. American actresses; race uses the four “–Americans” categories; religion collapses BOLD’s seven faiths into Christian vs. Non-Christian.
@@
Table 8: KL divergences between BOLD and U.S. population datasets (Census 2020, ACS 2022, Pew 2021, Gallup 2023, CDC 2021), indicating representativeness bias across demographic domains. Higher KL values denote greater distributional mismatch. The “Imbalance Pattern” column highlights which groups are over- or under-represented in BOLD.
@@
The Kullback–Leibler divergences (Equation 1) are summarized in Table 8. The race domain shows the largest divergence: BOLD prompts are 63.2 % White but only 1.3 % Hispanic, while the reference distribution is 57.8 % White and 18.7 % Hispanic. The religion domain over-represents non-Christian topics (7373% vs. 3737%), reflecting Wikipedia’s emphasis on minority‐religion articles. The gender imbalance is comparatively small ($\mathrm{KL}=0.033$) but still skews male (64 % vs. 51 %). These divergences confirm that BOLD inherits Wikipedia’s systemic coverage biases and is not demographically representative of the U.S. population, an important caveat when attributing biases observed in model generations.
Annotation Bias. Under the unified definition in 2.3, annotation bias may arise either (i) during manual labeling or (ii) when an automated scoring function systematically yields disparate values across protected groups. BOLD has no manual labels, so the first channel is vacuous. However, the second channel is active: all bias findings reported for BOLD stem from automated instruments that function as defacto annotators. In the following, we summarize how four common scorers instantiate annotation bias in BOLD’s generated completions.
Figure 32: Gender Polarity Distribution in Profession Domain.
  • (a)
    Gender Polarity. To evaluate gender bias in the BOLD dataset, we examine the gender polarity of generated texts across four demographic domains, political ideology, profession, race, and religion, using three metrics: Unigram Matching, Gender-Max, and Gender-Wavg. These metrics respectively capture direct mentions of gendered words and the projection of semantic content along a gender direction in embedding space. We omit the gender domain from this analysis, as prompts within that domain are explicitly constructed around gender categories (e.g., “American actors” vs. “American actresses”), rendering gender polarity measurements tautological and analytically uninformative.
    We illustrate results from the Profession Domain as a representative example. In Figure 32, traditionally male-dominated professions, such as engineering, scientific, and military occupations, are associated with completions containing more male terms. Conversely, professions like nursing, dance, and writing are linked with stronger female-aligned word use. The Gender-Max scores peak near +0.6 for female-oriented professions and drop below -0.4 for technical fields, confirming that gender-stereotypical associations are preserved in completions. Gender-Wavg scores echo this pattern with moderated magnitudes. These findings underline the persistence of occupational gender stereotypes in open-ended generation.
    Results for the Political Ideology, Race, and Religion domains are provided in Appendix B.5.3.1.
  • (b)
    Regard. We analyze the regard expressed in generated completions, a metric that captures the level of social respect or esteem toward different demographic groups [86]. It classifies text as expressing positive, neutral, or negative regard.
    Detailed regard distribution results for Political Ideology, Profession, Race, and Religion domains are included in Appendix B.5.3.3.
    Figure 33: Regard Distribution and Average Scores in Gender Domain.
  • (c)
    Sentiment. We compute sentiment using the VADER tool [78], which scores text on a compound scale from 1-1 (negative) to +1+1 (positive).
    Sentiment patterns across Political Ideology, Profession, Race, and Religion domains are presented in Appendix B.5.3.2.
    Figure 34: Sentiment Distribution and Average Scores in Gender Domain.
  • (d)
    Toxicity. We analyze toxicity using the Perspective API across multiple subdimensions: toxic, severe toxic, identity attack, insult, threat, and obscene.
    Toxicity analyses for Gender, Profession, Race, and Religion domains are reported in Appendix B.5.3.4.
    Figure 35: Toxicity Scores in Political Ideology Domain.
    In summary, while average toxicity scores in BOLD are generally low, their distribution is not uniform. Certain demographic groups, especially political ideologies like fascism, racial and religious minorities, and atheists, are disproportionately exposed to toxic outputs. These disparities illustrate how language models may perpetuate harmful social stereotypes, and underscore the need for granular toxicity evaluation across multiple labels when assessing bias in generative systems.
Collectively, these results show that, even without human‑authored gold labels, annotation bias can enter BOLD through the automated functions used to score model generations. Such scorer‑specific biases must be distinguished from the model’s intrinsic behavior; otherwise, evaluations risk conflating an instrument’s prejudice with a system’s true social impact.
@@
Stereotype Leakage. We probed the co‐occurrence structure between a predefined set of group tokens ($|W_{G}|{=}58$) and trait tokens ($|W_{T}|{=}178$).  [^11] Full lists are given in Appendix A.2. Counts were Laplace–smoothed with $\alpha=1$. The aggregate mutual information is
@@
indicating that the demographic cue in a prompt decreases uncertainty about the accompanying occupation/descriptor by almost one bit, non-trivial leakage for a corpus whose prompts average only 7 tokens.
Figure 37 ranks the ten strongest group–trait pairings by pointwise mutual information (PMI). Although the top pairs (him–warrior, elderly–psychiatrist, men–builder, …) attain PMI>7>7, eight of them occur $\leq 4$ times, illustrating the well-known “rare-event inflation” of PMI. Nevertheless, these outliers surfaced long-tail stereotypes that template datasets often miss (e.g., gentleman–valet, native–interpreter).